A bag contains 15 green 18 yellow and 16 orange balls one ball is randomly selected to the nearest percent what is the probability of the event
1
Expert's answer
2019-03-27T12:48:13-0400

The total number of balls in the bag

B=15+18+16=49

The probability that a random ball is green

Pgr=15/49=0.3061=30.61%

The probability that a random ball is yellow

Pye=18/49=0.3673=36.73%

The probability that a random ball is orange

Por=16/49=0.3265=32.65%

The probability that a random ball is green or yellow

Green balls+yellow balls=15+18=33

Pgr-ye=33/49=0.6735=67.35%

The probability that a random ball is green or orange

Green balls+orange balls=15+16=31

Pgr-or=31/49=0.6327=63.27%

The probability that a random ball is yellow or orange

Yellow balls+orange balls=18+16=34

Pye-or=34/49=0.6939=69.39%